OpenStudy (anonymous):

Four times a certain number is the same as the number increased by 78. Find the number. The number is ____. (Give only the value of the number as your answer.)

OpenStudy (shane_b):

4x=x+78 Solve for x

OpenStudy (anonymous):

4*x = y y = x+78 4*x = x+78 x= 26

OpenStudy (anonymous):

4x-x=78m so, 3x=78, x=78/3 x= 26

OpenStudy (shane_b):

Let's write it out another way just to be sure! \[\large 4x=x+78\]\[\large 4x-x=78\]\[\large 3x=78\]\[\large \frac{3x}{3}=\frac{78}{3}\]\[\large x=\frac{78}{3}\]\[\large \frac{78}{3}=26\]\[\large x=26\]